(please help me to solve this question)
solving equation zē-6z+9=4√(zē-6z-6) following roots are obtained
(A) 3+2√3, 3-2√3 (B) 5,1 (C) all the above (D) none

Answer by ewatrrr(24785)

cos(57) = cos(360-57) = cos(303)
.